OpenStudy (anonymous):

Solve for x: (2x − 14) = 4

OpenStudy (mathteacher1729):

What have you tried so far?

OpenStudy (anonymous):

Add 14 to both the sides..

OpenStudy (anonymous):

add 14 to both sides like this 2x - 14 +14 =4 +14 2x= 18 then you divide each side by 2 like this 2x/2= 18/2 x= 9 that's it